OpenStudy (anonymous):

-x+3y=6 x+3y=18 solve please?

OpenStudy (anonymous):

-x + 3y = 6 x + 3y = 18 ------------ 6y = 24 y = 4 ------------ x + 3y = 18 x + 12 = 18 x = 6

OpenStudy (anonymous):

thanks! I knewi could count on you! :D I don't understand this.

OpenStudy (anonymous):

It is just adding those two equations. -x and +x cancel out to zero. 3y + 3y becomes 6y. 6 + 18 becomes 24.